# Recursively-Constrained Partially Observable Markov Decision Processes

### v2

@@ 2.2 Pathological Behavior of Optimal Policies: Stochastic Self-Destruction
One consequence of BPO violation by the C-POMDP formulation in Problem 1 is that optimal policies may exhibit pathological behaviors during execution. Consider the POMDP in Figure 1. The optimal policy from $b_{0}$ first chooses action $a_{A}$ which incurs no cost. Suppose that $b_{1}$ is reached. The optimal policy plans to choose action $a_{A}$ and incur a cost of $8$ which violates the constraint, even though there is another action, $a_{B}$, that incurs a lower cost on expectation that satisfies the constraint. This means that, in $50\%$ of executions, when $b_{1}$ is reached, the agent intentionally violates the cost constraint to get higher expected rewards, even if a policy that does not violate the cost constraint exists. We term this pathological behavior stochastic self-destruction.
@@
This unintuitive behavior is chosen in the C-POMDP framework because the policy still satisfies the cost constraint at the initial belief state on expectation, so the policy is admissible. An optimal C-POMDP policy exploits the nature of the constraint in Eq. (6) to intentionally violate the cost constraint for some belief trajectories. A concrete manifestation of this phenomenon is in the stochasticity of the optimal policies for C-POMDPs. These policies randomize between deterministic policies that violate the expected cost threshold but obtain higher expected reward, and those that satisfy the cost threshold but obtain lower expected reward.
@@
Another consequence of violating BPO is a mismatch between optimal policies planned from a current belief state and optimal policies planned at future (successor) belief states. Consider the example in Figure 1 again. If re-planning is conducted at $b_{1}$, the re-planned optimal policy chooses $a_{B}$ instead of $a_{A}$. In fact, the policy that takes $a_{B}$ at $b_{0}$ achieves a higher expected reward than the policy that takes $a_{A}$ at $b_{0}$ and re-plans at future time steps. The violation of BPO can therefore lead to irrecoverable failures and poor performance of the closed-loop system at run time.
@@
It is important to note that the pathological behavior arises due to the C-POMDP problem formulation, and not the algorithms designed to solve C-POMDPs. Further, this issue cannot be addressed by simply restricting planning to deterministic policies since they also exhibit the pathological behavior as seen in the example in Figure 1.
@@ 3 Recursively-Constrained POMDPs
To mitigate the pathological behaviors and obtain an optimal substructure property, we aim to align optimal policies computed at a current belief with optimal policies computed at future (successor) beliefs. We propose a new problem formulation called Recursively-Constrained POMDP (RC-POMDP), which imposes additional recursively defined constraints on a policy.
@@
Let $h_{t}=\{a_{0},o_{0},\cdots,a_{t-1},o_{t-1}\}$ denote the history of the actions and observations up to time step $t$ but not including $t$; thus, $h_{0}=\emptyset$. The belief at time step $t$ can be computed as $b_{t}=P(s_{t}\mid b_{0},h_{t})$. For readability, we do not explicitly write $b_{0}$ in the notations, since all variables are conditioned on $b_{0}$. Then, the expected cumulative cost of the trajectories associated with history $h_{t}$ is given as:
@@
Since RC-POMDP policies are constrained based on history, it is not sufficient to use belief-based policies. Thus, we consider history-based policies in this work. A history-based policy maps a history $h_{t}$ to a probability over actions $\Delta(A)$. The RC-POMDP optimization problem is formalized below.
@@
Given a C-POMDP and an admissibility constraint $k\in\mathbb{N}\cup\{\infty\}$, compute optimal policy $\pi^{*}$ that is k-admissible, i.e.,
$\pi^{*}$ $=\arg\max_{\pi}V_{R}^{\pi}(b_{0})$ (12)
@@
In general POMDP problems, reasoning about cost is done based on expectation due to uncertainty. C-POMDPs bound the expected total cost of state trajectories; hence, belief trajectories with low expected costs can compensate for those with high expected costs. On the other hand, a worst-case constraint that does not use expectation never allows violations during execution, which may be overly conservative. RC-POMDPs provide a balance between the two; it bounds the expected total cost of all belief trajectories, only allowing cost violations during execution due to state uncertainty.
@@ 4 Theoretical Analysis of RC-POMDPs
RC-POMDP Problem 2 can be rewritten as:
$\begin{split}\pi^{*}&=\arg\max_{\pi}V_{R}^{\pi}(b_{0})\\ &\text{ s.t. }\quad V_{C}^{\pi}(b_{t})\leq d(h_{t})\qquad\forall t\in\{0,1,\dots,k-1\},\end{split}$ (16)
@@
Proof.
Consider an optimal (stochastic) policy $\pi^{*}$. Consider a $h_{t}$ reachable from $b_{0}$ by $\pi^{*}$, where $\pi^{*}$ has randomization (at least two actions have probabilities in $(0,1)$). Then, $\pi^{*}$ can be equivalently represented as a mixture over $n<|A|$ policies that deterministically selects a unique action at $h_{t}$ but is the same (stochastic) policy as $\pi^{*}$ everywhere else. That is, $\pi^{*}$ is a mixed policy over the set $\vec{\pi}^{*}=\{\pi_{1},\cdots,\pi_{n}\}$ of $n$ policies that have a deterministic action at $h_{t}$ and $\pi_{i}=\pi_{j}=\pi^{*}$ at every other history. Let $w_{i}$ represent the non-zero probability of choosing $\pi_{i}$ at $h_{t}$. Then, $V^{\vec{\pi}^{*}}_{R}(h_{t})=\sum_{i=1}^{n}w_{i}V^{\pi_{i}}_{R}(b_{t})$ and $V^{\vec{\pi}^{*}}_{C}(h_{t})=\sum_{i=1}^{n}w_{i}V^{\pi_{i}}_{C}(b_{t})$.
[…]
$[\forall h_{t+1}$ $,V_{C}^{\pi_{j}}(b_{t+1})\leq d(h_{t+1})]\implies V_{C}^{\pi_{j}}(b_{t})\leq d(h_{t}),$
seen by rearranging $V_{C}^{\pi_{j}}(b_{t})=C(b_{t},\pi_{j}(h_{t}))+\gamma\mathbb{E}[V_{C}^{\pi_{j}}(b_{t+1})]$ and Eq. (15). That is, Eq. (16) is violated at time step $t+1$, so $\vec{\pi}^{*}$ is inadmissible, which is again a contradiction. Hence, each $\pi_{i}\in\vec{\pi}^{*}$ is admissible.
Satisfaction of Bellman’s Principle of Optimality
Here, we show that RC-POMDPs satisfy BPO.
@@
The following theorem states that RC-POMDPs obey BPO.
@@
The Bellman backup operator for an RC-POMDP is given by, $\forall\bar{b}_{t}$,
$\mathbb{B}[V^{\pi}](\bar{b}_{t})=\max_{a\in A}\Big{[}Q_{R}^{\pi}((b_{t},d(h_{t})),a)\mid\\ Q^{\pi}_{C}((b_{t},d(h_{t})),a)\leq d(h_{t})\Big{]}.$ (17)
Further, assume an admissible policy exists for the RC-POMDP with admissibility constraint $k=\infty$. Then, optimal policy $\pi^{*}$ satisfies the BPO criterion
$\mathbb{B}[V^{\pi^{*}}](\bar{b}_{t})=V^{\pi^{*}}(\bar{b}_{t})\quad\forall\bar{b}_{t}\in\textsc{Reach}^{\pi^{*}}\!(\bar{b}_{0}),$
@@
Proof.
We first show that the evolution of $\bar{b}_{t}$ is Markovian, by showing $P(\bar{b}_{t+1}\mid b^{\prime}_{t},a_{t},o_{t},h_{t})=P(\bar{b}_{k+1}\mid\bar{b}_{t},a_{t},o_{t})$. See that
$P(\bar{b}_{t+1}\mid\bar{b}_{t},a_{t},o_{t},h_{t})$
=$=\begin{cases}P(b_{t+1}\mid b_{t},a,o_{t},h_{t})&\text{if }d(h_{t+1})=\gamma^{-1}(d(h_{t})-C(b_{t},a_{t}))\\ 0&\text{otherwise,}\end{cases}$
where $P(b_{t+1}|b_{t},a_{t},o_{t},h_{t})=P(b_{t+1}|b_{t},a_{t},o_{t})$
$=\sum_{s^{\prime}\in S}Z(s^{\prime},a_{t},o_{t})\sum_{s\in S}b_{t}(s)T(s,a_{t},s^{\prime}).$
Thus, $P(\bar{b}_{t+1}\mid\bar{b}_{t},a_{t},o_{t},h_{t})=P(\bar{b}_{t+1}\mid\bar{b}_{t},a_{t},o_{t})$, i.e., the evolution of $\bar{b}_{t}$ is Markovian.
Denote by $\Pi^{\prime}$ the set of deterministic admissible policies, i.e., for every $\pi\in\Pi^{\prime}$ and $t\in\mathbb{N}\cup\{\infty\}$, $V^{\pi}_{C}(b_{t})\leq d(h_{t})$, and let $A^{\prime}(b_{t})$ be the set of actions allowed by the policies in $\Pi^{\prime}$ at $b_{t}$. Then,
$\mathbb{B}[V_{R}^{\pi}](b_{t},d(h_{t}))$ $=\max_{a\in A^{\prime}}\Big{[}Q_{R}^{\pi}((b_{t},d(h_{t})),a)\mid\pi\in\Pi^{\prime}\Big{]},$
$=\max_{a\in A^{\prime}}\Big{[}Q_{R}^{\pi}(b_{t},a)\mid\pi\in\Pi^{\prime}\Big{]}.$
Since the equation has no dependence on $d(h_{t})$, we have
$\mathbb{B}[V_{R}^{\pi}](b_{t})=\max_{a\in A^{\prime}}\Big{[}Q_{R}^{\pi}(b_{t},a)\Big{]}\quad\text{s.t.}\quad\pi\in\Pi^{\prime}.$
Note this is the standard Bellman operator for the POMDP with only $\Pi^{\prime}$ policies obtained from RC-POMDP. Hence, the Bellman operator for the RC-POMDP is Eq. (17).
From the results of the Bellman operator for a POMDP (Hauskrecht 1997), $\mathbb{B}$ is a contraction mapping and has a single, unique fixed point, i.e., for an optimal $\pi^{*}\in\Pi^{\prime}$, $\mathbb{B}[V_{R}^{\pi^{*}}]=V_{R}^{\pi^{*}}$. Thus, $\pi^{*}$ satisfies the BPO criterion:
$\mathbb{B}[V_{R}^{\pi^{*}}](\bar{b}_{t})=V_{R}^{\pi^{*}}(\bar{b}_{t})\qquad\forall\bar{b}_{t}\in\textsc{Reach}^{\pi^{*}}\!(\bar{b}_{0}).\qed$
Theorems 2 and 3 show that (i) it is sufficient to search in the space of deterministic policies for an optimal one, and (ii) the optimal substructure property of RC-POMDPs can be exploited to employ standard dynamic programming methodology for an effective and computationally efficient approach (algorithm) to the RC-POMDP problem. These results also suggest that optimal policies for RC-POMDPs do not exhibit the same pathological behaviors as C-POMDPs.
@@ 5 Dynamic Programming for RC-POMDPs
We devise an algorithm as a first attempt to solve Problem 2 for RC-POMDPs with scalar cost and admissibility constraint $k=\infty$. We leave the multi-dimensional and finite $k$ case for future work. The algorithm is called Admissibility Recursively Constrained Search (ARCS). ARCS takes advantage of Theorems 2 and 3 to utilize point-based dynamic programming in the space of deterministic policies. ARCS uses techniques from POMDP tree search (Kurniawati, Hsu, and Lee 2008; Smith and Simmons 2005; Spaan and Vlassis 2005) to simultaneously search for admissible policies while improving bounds on reward-values.
ARCS is outlined in Alg. 1. It takes as input the RC-POMDP $\mathcal{M}$ and $\epsilon>0$, a target error between the computed policy and an optimal policy at the initial belief $b_{0}$. ARCS incrementally samples history-belief points as nodes in a policy tree $T$, and maintains the admissible cost, bounds on reward and cost values, and a lower bound on the admissible horizon from each node. At each iteration, a SAMPLE step picks a sequence of nodes starting from the root. Then, a Bellman BACKUP step is performed for each sampled node. Finally, a PRUNE step prunes sub-optimal nodes from the tree. These three steps are repeated until an admissible and $\epsilon$-optimal policy is found. Pseudocode for SAMPLE, BACKUP and PRUNE are provided in the appendix.
@@
1:  Initialize $\Gamma_{c_{min}}$, the cost-minimizing policy.
@@
3:  $\underline{V}_{R}\leftarrow\alpha_{r}^{T}b_{0},\overline{V}_{C}\leftarrow\alpha_{c}^{T}b_{0}$.
4:  Initialize $\overline{V}_{R}$ and $\underline{V}_{C}$ for $b_{0}$ with Fast Informed Bound.
5:  $V^{U}_{R}\leftarrow\overline{V}_{R}$, $V^{L}_{C}\leftarrow\underline{V}_{C}$.
6:  Initialize $k_{0}$ with Eq. (18)-(19).
7:  $T\leftarrow v_{0}=(b_{0},\mathcal{M}.\hat{c},k_{0},\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C},\emptyset,\emptyset,\emptyset,\emptyset)$.
8:  repeat
9:     $B_{sam}\leftarrow$ SAMPLE($\epsilon$).
10:     for all $v\in B_{sam}$ do
11:        BACKUP($v$).
12:        PRUNE($v$).
13:     end for
14:  until termination conditions are satisfied
15:  return $T,\Gamma_{c_{min}}$.
Policy Tree Representation
In POMDPs, $\alpha$-vectors (Sondik 1978) can be used for policy representation for a lower bound (upper bound) on reward-values (cost-values). However, for RC-POMDPs, the reward-value (cost-value) from an $\alpha$-vector policy may not be a lower bound (upper bound) on reward-value (cost-value) for an optimal admissible policy, so following an $\alpha$-vector policy may violate the cost constraints. Instead of a full $\alpha$-vector policy, we represent the policy with a policy tree $T$, and initialize lower bounds on reward-value (upper bounds on cost-value) for new sampled history-beliefs (leaf nodes) with a pre-computed minimum cost policy $\pi_{c_{min}}=\Gamma_{c_{min}}$ as a set of alpha-vector pairs $(\alpha_{r},\alpha_{c})$. The minimum cost policy is also used for history-beliefs that are not in $T$ during execution. $\pi_{c_{min}}$ can be computed efficiently using an unconstrained POMDP algorithm by defining a new reward function $R^{\prime}(s,a)=-C(s,a)$ and keeping track of the reward and cost $\alpha$-vectors. To initialize upper bounds of reward-values (lower bounds on cost-values) for new nodes, ARCS maintains an unconstrained upper reward-value $V^{U}_{R}$ and lower cost-value $V^{L}_{C}$ bounds, each initialized with the Fast Informed Bound technique (Hauskrecht 2000) for the unconstrained POMDP problem, separately using reward maximization and cost minimization.
@@
A node in $T$ is a tuple $v=(b,d,k,\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C},\overline{Q}_{R},\underline{Q}_{R},\overline{Q}_{C},\underline{Q}_{C})$, where $b$ is a belief state, $d$ is the history dependent admissible cost bound, $k$ is the admissible horizon guarantee, $\overline{V}_{R}$ and $\underline{V}_{R}$ are the upper and lower bound reward-values, $\overline{V}_{C}$ and $\underline{V}_{C}$ are the upper and lower bound cost-values, respectively. $\overline{Q}_{R},\underline{Q}_{R}$ represent the upper and lower bound on Q reward-value, and $\overline{Q}_{C},\underline{Q}_{C}$ represent upper and lower bounds on Q cost-value. The root of $T$ is the node $v_{0}$ with $b=b_{0}$ and $d=\hat{c}$, admissible horizon guarantee $k_{0}$.
@@
During SAMPLE and BACKUP, new nodes are initialized as follows. Belief $b^{\prime}$ is computed with Eq. (1), and $d^{\prime}$ is computed recursively with Eq. (15). $\underline{V}_{R}^{\prime}$ and $\overline{V}_{C}^{\prime}$ are initialized using the minimum-cost policy, $\overline{V}_{R}^{\prime}$ and $\underline{V}_{C}^{\prime}$ are initialized using the sawtooth approximation (Hauskrecht 2000) of $V^{U}_{R}$ and $V^{L}_{C}$, and $k^{\prime}$ is computed as a lower bound of the admissible horizon. Then, the node $v^{\prime}=(b^{\prime},d^{\prime},k,\cdot)$ is added to the tree as a successor to $v$.
Admissible Horizon Guarantee
@@
During search, it is computationally intractable to exhaustively search the possibly infinite space of an RC-POMDP to find admissible policies. Thus, we compute an admissible horizon guarantee $v.k$ for every node $v\in T$ as a lower bound on the admissible horizon of the policy. It is used to compute admissibility beyond the current search depth of the tree, and to improve search efficiency via pruning. To initialize the admissible horizon guarantee for a leaf node, we compute a lower bound on the admissible horizon $k$ of $\pi_{c_{min}}$, which is the minimum cost policy used for leaf nodes and future histories not in $T$.
@@
Let the maximum $1$-step cost $C_{max}$ that $\pi_{c_{min}}$ incurs at each time step across the entire belief space be $C_{max}=\max_{b\in B}C(b,\pi_{c_{min}}(b)).$ Then, for a node $v$ with $v.d<0$, $k=0$. For a leaf node $v$ with non-negative admissible cost $v.d\geq 0$, $\pi_{c_{min}}$ is at least $k$-admissible with
$k=\Big{\lfloor}\log\Big{(}1-\Big{(}\frac{v.d}{C_{max}}\Big{)}\cdot(1-\gamma)\Big{)}/\log(\gamma)\Big{\rfloor},$ (18)
and $\pi_{c_{min}}$ is admissible from history $h$ if
$(1-\gamma)^{-1}C_{max}\leq v.d\text{\;\;\; or \;\;\;}v.\overline{V}^{\pi_{c_{min}}}_{C}=0\leq v.d.$ (19)
@@
A proof is provided in the appendix. This lemma provides sufficient conditions for admissibility of computed policies. The only parameter is $C_{max}$, for which we compute an upper bound,
$C_{max}\leq V_{C,max}^{\pi_{c_{min}}}=\max_{b\in\Delta(S)}\min_{(\alpha_{r},\alpha_{c})\in\Gamma_{c_{min}}}\alpha_{c}^{T}b,$ (20)

### v3

@@ 2.1.1 Pathological Behavior : Stochastic Self-Destruction
In the above example, the optimal policy from $b_{0}$ first chooses action $a_{A}$. Suppose that $h_{1}$ is reached. The cost constraint at $b_{1}$ remains at $5$ since no cost has been incurred. However, the optimal C-POMDP policy chooses action $a_{A}$ and incurs a cost of $8$ which violates the constraint, even though there is another action, $a_{B}$, that incurs a lower expected cost that satisfies the constraint. Therefore, in $50\%$ of executions, when $h_{1}$ is reached, the agent intentionally violates the cost constraint to get higher expected rewards, even if a policy that satisfies the cost constraint exists. We term this pathological behavior stochastic self-destruction.
@@
This unintuitive behavior is mathematically correct in the C-POMDP framework because the policy still satisfies the constraint at the initial belief state on expectation. An optimal C-POMDP policy exploits the nature of the constraint in Eq. (6) to intentionally violate the cost constraint for some belief trajectories. A concrete manifestation of this phenomenon is in the stochasticity of the optimal policies for C-POMDPs. These policies randomize between deterministic policies that violate the expected cost threshold but obtain higher expected reward, and those that satisfy the cost threshold but obtain lower expected reward.
@@
Another consequence is a mismatch between optimal policies planned from a current time step and optimal policies planned at future time steps. In the example in Figure 1, if re-planning is conducted at $b_{1}$, the re-planned optimal policy selects $a_{B}$ instead of $a_{A}$. In fact, the policy that initially takes $a_{B}$ at $b_{0}$ achieves a higher expected reward than the original policy that takes $a_{A}$ at $b_{0}$ and re-plans at future time steps. This phenomenon can therefore lead to poor performance of the closed-loop system during execution.
@@
We remark that the pathological behavior arises due to the C-POMDP problem formulation, and not the algorithms designed to solve C-POMDPs. Further, this issue cannot be addressed by simply restricting solutions to deterministic policies since they also exhibit the pathological behavior, as seen in the example in Figure 1.
@@ 3 Recursively-Constrained POMDPs
To mitigate the pathological behaviors and obtain a (policy-independent) optimal substructure property, we aim to align optimal policies computed at a current belief with optimal policies computed at future (successor) beliefs. We propose a new problem formulation called Recursively-Constrained POMDP (RC-POMDP), which imposes additional recursively defined constraints on a policy.
@@
The expected cumulative cost of the trajectories associated with history $h_{t}$ is given as:
@@
Since RC-POMDP policies are constrained based on history, it is not sufficient to directly use belief-based policies. Thus, we consider history-based policies in this work. A history-based policy maps a history $h_{t}$ to a probability over actions $\Delta(A)$.
The RC-POMDP optimization problem is formalized below.
@@
Given a C-POMDP and an admissibility constraint $k\in\mathbb{N}\cup\{\infty\}$, compute optimal policy $\pi^{*}$ that is k-admissible, i.e., $\forall h_{t}$,
$\pi^{*}(h_{t})=\arg\max_{\pi}V_{R}^{\pi}(h_{t})$ (10)
@@
In POMDPs, reasoning about cost is done on expectation due to state uncertainty. C-POMDPs bound the expected total cost of state trajectories, enabling belief trajectories with low expected costs to compensate for those with high expected costs. Conversely, a worst-case constraint formulation of the problem, which never allows any violations during execution, may be overly conservative. RC-POMDPs strike a balance between the two; it bounds the expected total cost for all belief trajectories, only allowing cost violations during execution due to state uncertainty.
@@ 4 Theoretical Analysis of RC-POMDPs
Problem 2 can be rewritten as:
$\begin{split}\pi^{*}&=\arg\max_{\pi}V_{R}^{\pi}(b_{0})\\ &\text{ s.t. }\quad V_{C}^{\pi}(b_{t})\leq d(h_{t})\quad\forall t\in\{0,1,\dots,k-1\},\end{split}$ (14)
@@
A proof is provided in the Appendix. The main intuition is that we can always construct an optimal deterministic policy from an optimal stochastic policy. That is, at every history in which the policy has stochasticity, we can construct a new admissible policy that achieves the same reward-value while remaining admissible by deterministically choosing one of the stochastic actions at that history. We obtain a deterministic optimal policy by inductively performing this determinization at all reachable histories.
Satisfaction of Bellman’s Principle of Optimality   Here, we show that RC-POMDPs satisfy BPO with a policy-independent optimal substructure.
@@
We first see that the evolution of $\bar{b}_{t}$ is Markovian, i.e.,
$P(\bar{b}_{t+1}\mid\bar{b}_{t},a_{t},o_{t},h_{t})$
=$=\begin{cases}P(b_{t+1}\mid b_{t},a,o_{t},h_{t})&\text{if }d(h_{t+1})=\frac{(d(h_{t})-C(b_{t},a_{t}))}{\gamma}\\ 0&\text{otherwise,}\end{cases}$
thus, $P(\bar{b}_{t+1}\mid\bar{b}_{t},a_{t},o_{t},h_{t})=P(\bar{b}_{t+1}\mid\bar{b}_{t},a_{t},o_{t})$.
Here, we use the policy iteration version of Bellman equation, but a similar argument can be made for value iteration.
@@
Fix $\pi$. Let $V^{\pi}=(V_{R}^{\pi},V_{C}^{\pi})$ be reward- and cost-value function for $\pi$. The Bellman operator $\mathbb{B}$ for policy $\pi$ for an RC-POMDP is given by, $\forall\bar{b}_{t}$,
$\mathbf{a}=\operatorname*{arg\,max}_{a\in A}\Big{[}Q_{R}^{\pi}(\bar{b}_{t},a)\mid Q^{\pi}_{C}(\bar{b}_{t},a)\leq d(h_{t})\Big{]}$ (15)
$\mathbb{B}[V^{\pi}](\bar{b}_{t})\triangleq\begin{cases}\big{(}Q_{R}^{\pi}(\bar{b}_{t},a),Q_{C}^{\pi}(\bar{b}_{t},a)\big{)},\,a\in\mathbf{a}&\text{ if }\mathbf{a}\neq\emptyset,\\ \big{(}V_{R}^{\pi}(\bar{b}_{t}),(\infty,\ldots,\infty)\big{)}&\text{ if }\mathbf{a}=\emptyset,\end{cases}$
Assume an admissible policy exists for the RC-POMDP with admissibility constraint $k=\infty$. Let $V^{\pi^{*}}=(V_{R}^{\pi^{*}},V_{C}^{\pi^{*}})$ be the values for an optimal admissible policy $\pi^{*}$, and we obtain a new policy $\pi^{\prime}$ with $(V_{R}^{\pi^{\prime}},V_{C}^{\pi^{\prime}})=\mathbb{B}[V^{\pi^{*}}]$. $\pi^{*}$ satisfies the BPO criterion of an admissible optimal policy:
$V_{R}^{\pi^{\prime}}(\bar{b}_{t})$ $=V_{R}^{\pi^{*}}(\bar{b}_{t})\qquad\forall\bar{b}_{t},$ (16)
$V_{C}^{\pi^{\prime}}(b_{t})$ $\leq d(h_{t})\qquad\quad\,\forall\bar{b}_{t}\in\textsc{Reach}^{\pi^{\prime}}\!(\bar{b}_{0}),$ (17)
@@
This theorem shows that an optimal policy remains admissible and optimal w.r.t rewards after applying $\mathbb{B}$ on a policy independent value function $V$. Note that $V_{C}^{*}$ is not unique as there may be multiple optimal cost-value functions for an optimal $V_{R}^{*}$. Next, we show that $\mathbb{B}$ is a contraction over reward-values for a suitably initialized value function, which is one that defines the space of admissible policies.
Theorem 3.
For each $\bar{b}_{t}$, define $\Phi(\bar{b}_{t})$ as the set of admissible policies from $\bar{b}_{t}$:
$\Phi(\bar{b}_{t})=\{\pi\;|\;V^{\pi}_{C}(b_{\tau})\leq d(h_{\tau})\;\;\forall\tau\geq t\}.$ (18)
$V^{\pi^{0}}$ is a well behaved initial value function if the following holds for all $\bar{b}_{t}$. If $\Phi(\bar{b}_{t})=\emptyset$, $V^{\pi^{0}}_{C}(\bar{b}_{t})=(\infty,\ldots,\infty)$. If $\Phi(\bar{b}_{t})\neq\emptyset$, $V_{C}^{\pi^{0}}(\bar{b}_{t})\leq d(h_{t})$.
Suppose that $V^{\pi^{0}}$ is well behaved, then $\mathbb{B}^{n}[(V_{R}^{\pi^{0}},V_{C}^{\pi^{0}})]_{x}\rightarrow(V_{R}^{\pi^{*}},V_{C}^{\pi^{n}})$ as $n\rightarrow\infty$. That is, starting from $\pi^{0}$, $\mathbb{B}$ is a contraction on $V_{R}$ and $V_{R}^{\pi^{*}}$ is a unique fixed point.
Proofs of all results are provided in the Appendix. Theorems 1-3 show that it is sufficient to search in the space of deterministic policies for an optimal one, and the policy-independent optimal substructure of RC-POMDPs can be exploited to employ dynamic programming for an effective and computationally efficient algorithm for RC-POMDPs. Further, Theorem 3 shows that determining policy admissibility is essential for effective dynamic programming. These results also indicate that optimal policies for RC-POMDPs do not exhibit the same pathological behaviors as C-POMDPs.
@@ 5 Dynamic Programming for RC-POMDPs
With the theoretical foundation above, we devise a first attempt at an algorithm that approximately solves Problem 2 with scalar cost and admissibility constraint $k=\infty$. We leave the multi-dimensional and finite $k$ cases for future work. The algorithm is called Admissibility Recursively Constrained Search (ARCS). ARCS takes advantage of the Markovian property of the belief-admissible cost formulation in Proposition 2, and Theorems 1-3 to utilize point-based dynamic programming in the space of deterministic and admissible policies, building on unconstrained POMDP methods [Shani et al., 2013].
ARCS is outlined in Algorithm 1. It takes as input the RC-POMDP $\mathcal{M}$ and $\epsilon>0$, a target error between the computed policy and an optimal policy at $b_{0}$. ARCS explores the search space by incrementally sampling points in the history space. These points form nodes in a policy tree $T$. At each iteration, a SAMPLE step expands a sequence of points starting from the root. Then, a Bellman BACKUP step is performed for each sampled node. Finally, a PRUNE step removes sub-optimal nodes. These three steps are repeated until an admissible $\epsilon$-optimal policy is found. Pseudocode for SAMPLE, BACKUP and PRUNE are provided in the appendix.
@@
1:  Initialize cost-minimizing policy $\hat{\pi}_{c}^{min}=\Gamma_{c}^{min}$
@@
3:  $\underline{V}_{R}\leftarrow\alpha_{r}^{T}b_{0},\overline{V}_{C}\leftarrow\alpha_{c}^{T}b_{0}$
4:  Initialize $\overline{V}_{R}$ and $\underline{V}_{C}$ for $b_{0}$ with FIB
5:  Initialize $k_{0}$ with Eq. (19)-(20)
6:  $T\leftarrow v_{0}=(b_{0},\hat{c},k_{0},\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C},\emptyset,\emptyset,\emptyset,\emptyset)$
7:  repeat
8:     $B_{sam}\leftarrow$ SAMPLE($\epsilon$).
9:     for all $v\in B_{sam}$ do
10:        BACKUP($v$)
11:     end for
12:     PRUNE($B_{sam}$)
13:  until termination conditions are satisfied
14:  return $T,\Gamma_{c_{min}}$
@@
Policy Tree Representation   We represent the policy with a policy tree $T$. A node in $T$ is a tuple $v=(b,d,k,\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C},\overline{Q}_{R},\underline{Q}_{R},\overline{Q}_{C},\underline{Q}_{C})$, where $b$ is a belief, $d$ is a history-dependent admissible cost bound, $k$ is a lower bound on admissible horizon, $\overline{V}_{R}$ and $\underline{V}_{R}$ are the two-sided bounds on reward-values, $\overline{V}_{C}$ and $\underline{V}_{C}$ are the two-sided cost-value bounds, $\overline{Q}_{R},\underline{Q}_{R}$ represent two-sided bound on $Q$ reward-value, and $\underline{Q}_{C},\overline{Q}_{C}$ represent the two-sided bounds on $Q$ cost-value. The root of $T$ is the node $v_{0}$ with $b=b_{0}$, $d=\hat{c}$, and admissible horizon lower bound $k_{0}$.
From Theorem 3, a key aspect of effective dynamic programming for RC-POMDPs is computing admissible policies. This can be approximated by minimizing $V_{C}(\hat{b}_{t})$. As a pre-processing step, we first approximate a minimum cost-value policy $\pi_{c}^{min}=\arg\inf_{\pi}V^{\pi}_{C}$. An arbitrarily tight under-approximation (upper bound) $\hat{\pi}_{c}^{min}$ as a set of $|S|$-dimensional hyperplanes, called $\alpha$-vectors, can be computed efficiently with a POMDP algorithm Hauskrecht [2000]. The reward-values obtained by $\hat{\pi}_{c}^{min}$ is also a lower bound on the optimal reward-value. Thus, $\hat{\pi}_{c}^{min}$ is represented by a set of $\alpha$-vector pairs $(\alpha_{r},\alpha_{c})\in\Gamma_{C}^{min}$. $\hat{\pi}_{c}^{min}$ is used to initialize our policy, and is used from leaf nodes of $T$.
@@
To initialize a new node, belief $b^{\prime}$ is computed with Eq. (1), and $d^{\prime}$ is computed recursively with Eq. (13). We initialize $\underline{V}_{R}$ and $\overline{V}_{C}$ with $\hat{\pi}_{c}^{min}$, and initialize $\underline{V}_{C}$ and $\overline{V}_{R}$ independently using the Fast Informed Bound (FIB) Hauskrecht [2000], and $k^{\prime}$ is a lower bound of the admissible horizon.
@@
Admissible Horizon Lower Bound   It is computationally intractable to exhaustively search the possibly infinite policy space. Thus, we maintain a lower bound on the admissible horizon of the policy for every node. It is used to compute admissibility beyond the current search depth of the tree, and to improve search efficiency via pruning. To initialize the admissible horizon guarantee of a leaf node, we compute a lower bound on the admissible horizon $k$ when using $\hat{\pi}_{c_{min}}$.
@@
Let the maximum $1$-step cost $C_{max}$ that $\hat{\pi}_{c}^{min}$ incurs at each time step across the entire belief space be $C_{max}=\max_{b\in B}C(b,\hat{\pi}_{c}^{min}(b)).$ Then, for a node $v$, if $v.d<0$, then $k=0$. For a leaf node $v$ with $v.d\geq 0$, $\hat{\pi}_{c}^{min}$ is at least $k$-admissible with
$k=\big{\lfloor}\log\big{(}1-({v.d}/{C_{max}})\cdot(1-\gamma)\big{)}/\log(\gamma)\big{\rfloor},$ (19)
and $\hat{\pi}_{c}^{min}$ is admissible from history $h$ if
${C_{max}}/{(1-\gamma)}\leq v.d\text{\;\; or \;\;}v.\overline{V}^{\hat{\pi}_{c}^{min}}_{C}=0\leq v.d.$ (20)
@@
A proof is provided in the appendix. This lemma provides sufficient conditions for admissibility of computed policies. We compute an upper bound on the parameter $C_{max}$,
$C_{max}\leq V_{C,max}^{\hat{\pi}_{c}^{min}}=\max_{b\in\Delta(S)}\min_{(\alpha_{r},\alpha_{c})\in\Gamma_{c_{min}}}\alpha_{c}^{T}b,$ (21)